# Bootstrapping the Abelian Lattice Gauge Theories

### v1

@@ 6. Conclusion and Discussion
    The loop equations and positivity conditions can numerically pin down the 2D $U(1)$ lattice gauge theory.
  • We bootstrapped two classical Abelian lattice gauge theories: the $U(1)$ and $\mathbb{Z}_{2}$ lattice gauge theories on the 3D cubic lattice. We obtained two-sided bounds on the plaquette average. The two-sided bounds quickly converge in the strong or weak coupling limit with high precision. We compared our numerical results with the perturbative expansions in the strong or weak coupling limit, as well as the Monte Carlo data, and find perfect agreement. The numerical precision of the two-sided bounds is less impressive for certain intermediate coupling coefficients, which we expect can be significantly improved by bootstrapping Wilson loops with higher $L_{\textrm{max}}$.
@@
Perhaps the most important problem is to extend the 2D lattice bootstrap with high $L_{\textrm{max}}$ to higher dimensions. The 2D $U(1)$ results are encouraging for bootstrapping higher dimensional lattice gauge theories. It is tempting to believe that for higher dimensional lattice gauge theories, with sufficient large $L_{\textrm{max}}$ the bootstrap bounds may also converge to solutions with high precision. While this is hard to be verified in practise, as the bootstrap implementations quickly become too complicated to study. The number of Wilson loops increases exponentially with higher $L_{\textrm{max}}$, but empirically only small part of them have notable contributions on the bootstrap bounds. It would be important to find systematical selection rules for the Wilson loops to improve the numerical efficiency. This problem gets more subtle as the Wilson loops are connected with each other through loop equations.
@@
It is interesting that with a large coupling coefficient $\lambda$, the 3D two-sided bootstrap bounds merge and coincide with the 2D theories. It has been predicted using the Griffiths’ inequalities long ago that the 2D theories provide lower bounds for the Wilson loop averages in higher dimensions. While our bootstrap bounds suggest the Griffiths lower bound is actually optimal with large couplings. Physically this may relate to the fact that with strong coupling the gauge interactions are confined in a small region so are insensitive to the global geometry of the lattice. Nevertheless, it is interesting that the result can be obtained just from positivity/convexity reasons. It would be interesting to explore the constraints from the positivity and convexity further and study the universal behaviors of the confinement. Another interesting question is whether there are similar properties for the Yang-Mills gauge theories. The convex cones in the Yang-Mills lattice gauge theories are quite subtle. One may firstly gain evidence on this problem by comparing the bootstrap bounds of the same gauge group but in different dimensions [34, 35, 60], and check if the higher dimensional bounds are strictly above while converging to the 2D solutions with large coupling.
We presented the primary bootstrap bounds which can be used to estimate the string tension and glueball mass spectrum. Interesting data on these parameters could be obtained given the precision of the bootstrap bounds can be improved by one order. Motivated by the 2D bootstrap results on the precision $\delta w_{1}$ and the maximum length of the Wilson loops $L_{\textrm{max}}$, we expect the precision of the 3D bootstrap bounds could be significantly improved by increasing $L_{\textrm{max}}$ from our current value $16$ to $24$. We leave these explorations for future work.
@@ Acknowledgements
We would like to thank Zechuan Zheng for very helpful discussions and communications. This research was supported by the Startup Funding 4007022314 of the Southeast University, and the National Natural Science Foundation of China funding No. 12375061.

### v2

@@ 6. Conclusion and Discussion
    With large $L_{\textrm{max}}$, the loop equations and positivity conditions can numerically pin down the 2D $U(1)$ lattice gauge theory.
  • We bootstrapped two classical Abelian lattice gauge theories: the $U(1)$ and $\mathbb{Z}_{2}$ lattice gauge theories on the 3D cubic lattice. We obtained two-sided bounds on the plaquette average. The two-sided bounds quickly converge in the strong or weak coupling limit with high precision. We compared our numerical results with the perturbative expansions in the strong or weak coupling limit, as well as the Monte Carlo data and found excellent agreements. The numerical precision of the two-sided bounds is less impressive for certain intermediate coupling coefficients, which we expect can be significantly improved by bootstrapping Wilson loops with higher $L_{\textrm{max}}$.
@@
Perhaps the most important problem is to extend the 2D lattice bootstrap with high $L_{\textrm{max}}$ to higher dimensions. The 2D $U(1)$ results are encouraging for bootstrapping higher dimensional lattice gauge theories. It is tempting to believe that for higher dimensional lattice gauge theories, with sufficiently large $L_{\textrm{max}}$ the bootstrap bounds may also converge to solutions with high precision. Convergence of the bootstrap bounds on the lattice Ising model has been studied in [84]. While this is hard to be verified in practise, as the bootstrap implementations quickly become too complicated to study. The number of Wilson loops increases exponentially with higher $L_{\textrm{max}}$, but empirically only small part of them have notable contributions on the bootstrap bounds. It would be important to find systematical selection rules for the Wilson loops to improve the numerical efficiency. This problem gets more subtle as the Wilson loops are entangled with each other through loop equations.
@@
It is interesting that with a large coupling coefficient $\lambda$, the 3D two-sided bootstrap bounds merge and coincide with the 2D theories. It has been predicted using the Griffiths’ inequalities long ago that the 2D theories provide lower bounds for the Wilson loop averages in higher dimensions [72]. While our bootstrap bounds suggest the Griffiths lower bound is actually optimal with large couplings. Physically this may relate to the fact that with strong coupling the gauge interactions are confined in a small region so are insensitive to the global geometry of the lattice. Nevertheless, it is remarkable that the result can be obtained just from positivity/convexity reasons. It would be interesting to explore the constraints from the positivity and convexity further and study the universal behaviors of the confinement. Another interesting question is whether there are similar properties for the Yang-Mills gauge theories. The persumed convex cones in the Yang-Mills lattice gauge theories are quite subtle. One may firstly gain evidence on this problem by comparing the bootstrap bounds of the same gauge group but in different dimensions [36, 37, 64], and check if the higher dimensional bounds are strictly above while converging to the 2D solutions with large coupling.
We presented the primary bootstrap bounds which can be used to estimate the string tension and glueball mass spectrum. Interesting data on these parameters could be obtained given the precision of the bootstrap bounds can be improved by one order better. Motivated by the 2D bootstrap results on the precision $\delta w_{1}$ and the maximum length of the Wilson loops $L_{\textrm{max}}$, we expect the precision of the 3D bootstrap bounds could be significantly improved by increasing $L_{\textrm{max}}$ from our current value $16$ to $24$. We leave these explorations for future work.
@@ Acknowledgements
We would like to thank Zechuan Zheng for very helpful discussions and communications. We are grateful to Jiaxin Qiao for the comments on the first version of this paper. This research was supported by the Startup Funding 4007022314 of the Southeast University, and the National Natural Science Foundation of China funding No. 12375061.